This question is on a quadratics practice, but I'm stumped as to how quadratics can even be applied here.
Mr. Gordon bought 100 feet of fencing material to build a kennel for his dog. He is using one side of his garage for one side of the kennel. Write a quadratic function to represent the area of the kennel.

Answer by ikleyn(52781)

.

If W is the dimension of the kennel in direction perpendicular to wall, then the dimension parallel to wall is (100-2W) feet.


The area of the kennel is then  area = w*(100-W) = -W^2 + 100W.


It is that quadratic function, the question is asking for.